OpenStudy (anonymous):

Which of the quadratics has a graph with x-intercepts of 5 and -1? A) y = x2 - 4x - 5 B) y = x2 - 5x - 4 C) y = x2 + 4x - 5 D) y = 2x2 - 4x - 5

OpenStudy (reemii):

the x-intercepts are the roots right?

OpenStudy (reemii):

alll polynomials that have roots 5 and -1 are of the form \(a(x-5)(x+1) = a(x^2-4x-5)\). Find the one in your options that looks like that. (\(a\) may be equal to 1, -1, 2, .. something..)

OpenStudy (anonymous):

check A y = x2 - 4x - 5